INAF queue observing, 2017 Sep 26-27

Observer: A. Rossi, R. Carini
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCI1, LUCI2

Summary

At the sunset the sky is partially cloudy. At 11:10 (UT) the sky is totally covered, we stop the acquisition and close the dome

Night duration (12\xB0 el twilight) 2:00am- 12:20pm UT = 10 h 20m

Observing time: * *7.6hrs (73%)

Weather time loss: * *0.7h (7%)

technical time loss * *2h (20%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
bias   LBC      
dark   LUCI1,2      
telluric A12 HIP106817 LUCI1,2 HKspec 2m no other telluric of Piconcelli visible
Piconcelli A11 1633+3629 LUCI1,2 HKspec 0.3 completed
new telluric A11 HIP87643 LUCI1,2 HKspec   much closer to A11, GOOD only for LUCI2
Piconcelli A12 2238-0808 LUCI1,2 HKspec 0.8 completed
telluric A12 HIP106817 LUCI1,2 2m HKspec 0.2  
Rossi GRB170714A too LUCI1,2 H 30+30min completed
Cusano NW1 LBC B (blue channel), V (red channel) 14min the second ob is not good
Cusano NW2 LBC B (blue channel), V (red channel) 14min the second ob is not good
Cusano NW3 LBC B (blue channel), V (red channel) 7min ob 7
Dotto 2004TK10 BVg (blue), BzRI (red) 4m 30s repeated twice. sky cloudy and seeing bad

Details

darks Piconcelli A11 Dit =180, read-out mer.
  • luci1.20170927.0101-0105 not good, there are some persistences
  • luci2.20170927.0004-0008
  • in LIR mode the darks of LUCI1 are good, we try again in MER mode. the persistences are present again

02:15 start with telluric for Piconcelli A11
  • we choose the A12 telluric (HIP106817) because the one of A11 (HIP71172) is too low right now
  • GO TargetS _LS_200H+K_HKspec_tell_bino.xml
  • luci1.0110-0112 acquisition
  • luci2.0010-0011 acquistion
  • the star seems double, but the orientation of the stars are different in the two instruments, so it is probably a ghost, we take one other image.
  • luci1.0113, luci2.0012 now the star is single, we can continue the the acquisition luci1.0114,-luci2.0013-0014
  • luci1.0015-0016, luci2.0015-0016 science

02:53 start the target A11
  • during the acquisition of the source +slit image, LUCI stuck.
  • we take the image by hand, we verify that the source is in the slit and continue the script with the science, but during commit message of ERROR un LUCI2, we continue with LUCI1, after the first spectra (luci1.0121) also LUCI1 stuck.
  • 03:25 load again the script, we skip the acquisition part, now LUCi1 and LUCI2 work.
  • luci1.0122-0124 science
  • luci2.0021-0023 science
  • airmass about 1.49, seeing 0.8''

03:35 telluric for A11: we created a new script for the telluric HIP87643 : this telluric is better because it is much closer to A11 than the telluric for A12
  • 03:40 LUCI2 stuck
  • 03:52 load again the script and start
  • 04:00 acquisition LUCI1 failed, we do it by hand
  • we understand that LUCI1 has problems with the calculation of the offset
  • we calculate manually the offset and send them, we take the image and verify that the star is inside the slit (luci1.0134), but when we take the data the spectra there isn't (luci1.0135-0138), the problem could be linked with the guide probing.
  • LUCI2 works: acquisition: luci2.00024-0026, science: luci2.0027-0030
  • frozen RTD of LUCI1

04:46 slew to target A12 of Piconcelli Program
  • luci1.0139-0.143, luci2.0031-0034 acquisition
  • luci1.0144-0149, luci2.0035-0040 science
  • seeing 0.7'''

05:45 we slew to the telluric for A12 again (HIP106817 )
  • 05:55 LUCI1 and LUC2 stuck, we start again
  • acquisition luci1.0150-0152, luci2.0041-0043
  • science luci1.0153-0154, luci2.0043-0044
  • seeing 0.7'''

06:10 we slew to Nisini program, Field1
  • acquisition telluric Hip16652
  • 06:19 LUCi2 has not guide star, but LUCI1 has it
  • there are a lot of clouds
  • 06:33 there are too much clouds, the telescope can not guide, we stop
  • 06:50 start again but LUCI stuck
  • load again the script and re-start the acquisition
  • 07:13 error again copointing limits after send the offset
  • 07:32 we can not resolve the problem, we try with another target

07:33 slew to Rossi program
  • luci1.0161-0190, luci2.0052-0081
  • seeing is variable between 1.1'' and 1.5''
  • 08:13 the sky is clear and the seeing is more stable around 1.1''-1.2''

08:25 move to LBC

09:02 slew to Cusano NW-1 7th set of ob with thin clouds
  • dimm seeing 1"
  • dohybrid
  • copointing
  • start NW-1 ob7 at 9:18, seeing 1.2 on guider" at 09:20
  • dimm seeing 1" at 9:25
  • seeing 0.8" at 9:30 on LBCred image
  • start NW-2 ob7
  • seeing 0.8" at 9:35 on LBCred image
  • start NW-3 ob7
  • dofpia because dense cirrus just arrived frown, sad smile
  • start NW-1 ob8 at 9:55
  • start NW-2 ob8 at 10:05 but thick cirrus

10:14 slew to Rossi GRB170714A LBC
  • dofpia
  • start OB at 10:20
  • seeing 1" at 10:20

10:40 slew to Dotto 2004TK10 * dofpia * copointing * start non sideral tracking at 10:58 * seeing 1.8. * 11:08 seeing about 1.4''. * we have run the script twice because the sky is very cloudy and the seeing bad.

11:09 photometric standard. the sky is very cloudy
  • 11:10 the sky is totally covered, we stop the acquisition.

*calibration luci
  • flat for Rossi target luci1.0191-0200, luci2.0082-0091 (H filter)no good
  • flat and arc for Piconcelli program ==luci1.0201-0226, luci2.0092-==no good
  • flat Piconcelli program luci1.0227-0231, luci2.0113-0117
  • arc Piconcelli
  • neon luci1.0232-0233, luci2.0118-0119 ,
  • argon luci.0234-0235,luci2.0120-0121
  • xenon luci1.0236-0237, luci2.0122-0123
  • flat Rossi not good

* flat for Jochen
  • Fe II bright and Br gamma luci1.0238-..., luci2.0124-... Useless because N30 field stop not alligned

-- %USERSIG{RobertaCarini - 2017-09-27}%

Comments

 
Topic revision: r16 - 17 May 2018, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback